Shift Summary 10/05/99 Owl

Jason LaDue (jladue@cuhep.Colorado.EDU)
Tue, 5 Oct 1999 06:05:28 -0600 (MDT)

Shift Summary for owl shift, 10/05/99

On Shift: E. Turner, J. LaDue

Hours of beam available at all intensities: 0.00
Hours of beam available at nominal intensity: 0.00
Hours of E799 data (special runs not included): 0.00
Hours of special runs (laser scans, muon runs...): 0.00

Accounting of lost time:
Liquid Helium compressor problems with the Tevatron.

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:
Access to try and reset the Sector 3 LV supply.

========================================================================
Overview of shift:
No Beam for at least 24 Hours because of liquid helium
compressor problems. The couch again tempts another victim.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14802 00:00 New shift takes over cosmics running.

04:07 Val speaks to us, quick look at the online plots
reveals what looks to be a dead sector. SDAQ
window verifies the shift crews hypothesis.
Sector 3 is dead. Page Julie.

04:35 Go down on access to try to reset the power
supply. The -2.5 supply had a red OC light.
Attempts to reset the power supply fail. Julie
says that we might as well hold off until morning.
Stop run and wait.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Page BobT at a reasonable hour (ie 8:00AM) and inform him about the dead
LV supply.